Beaker A contains 1 liter which is 15 percent oil and the rest is vinegar, thoroughly mixed up. Beaker B contains 2 liters which is 40 percent oil and the rest vinegar, completely mixed up. Half of the contents of B are poured into A, then completely mixed up. How much oil (in liters)should now be added to A to produce a mixture which is 60 percent oil?
i dont even know where to begin!! any help would be really appreciated :)

Answer by stanbon(75887)

Beaker A contains 1 liter which is 15 percent oil and the rest is vinegar, thoroughly mixed up.
Oil = 0.15*1 = 0.15 liter ; total = 1 liter
---------------------------------
Beaker B contains 2 liters which is 40 percent oil and the rest vinegar, completely mixed up.
Oil = 0.40*2 = 0.80 liter ; total = 2 liter
---------------------------------
Half of the contents of B are poured into A, then completely mixed up.
That half contains 0.40 liter of oil ; total = 2 liter
---------------------------------
How much oil (in liters)should now be added to A to produce a mixture which is 60 percent oil?
---
Equation:
oil + oil = oil
0.40 + x = 0.60(2 + x)
Multiply thru by 100 to get:
40 + 100x = 60*2 + 60x
40x = 80
x = 2 liters (amount of oil that must be added)
======================================================
Cheers,
Stan H.